OpenStudy (anonymous):

Help me find out where I went wrong-which of the following have the same result? f(2)=3x+2, f^-1(3)=(2x-7)/3, 2y+14=4y-2 I got 8 for all of them.

OpenStudy (anonymous):

f(2)=3.2+2=6+2=8 f^-1(3)=(2.3-7)/3=6-7/3=-1/3 4y-2y-16=0 2y=16 y=8 so the first and last one are equal to 8

OpenStudy (anonymous):

Thank you, I see where I went wrong

OpenStudy (anonymous):

any time :)